# Giant High-order Nonlinear and Nonreciprocal Electrical Transports Induced by Valley Flipping in Bernal Bilayer Graphene

### v1

@@ V Discussion
And the odd and even parities of the normal current $j_{N,\perp}$ and Hall current $j_{H}$ under mirror operation are clearly shown in Fig. 7 by the last two columns.
Another concerning problem is spatial domains where the holes populate different mini-valleys, which make the system nonuniform. However, as shown in our paper, the valley polarizations in different domains could be aligned by a large enough electrical field or equivalently current flow. And the discussion of the uniform case in our paper is sufficient to describe realistic experiments.
Acknowledgements.
We thank Prof. Guangyu Zhang for their inspiring experimental observations and helpful discussions. This work was fully supported by a fellowship award from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. HKUST SRFS2324-6S01)
Supplemental Materials: Giant High-order Nonlinear and Nonreciprocal Electrical Transports Induced by Valley Flipping in Bernal Bilayer Graphene
I The Exchange Energy Functional
To minimize $E_{tot}(n_{1},n_{2},n_{3})$ in the main text, the exchange energy functional is further approximated by its series expansions with respect to $n_{\alpha=1,2,3}$. For convenience, we define the intra-valley exchange energy $E_{ex}^{i}$ and inter-valley exchange $E_{ex}^{x}$ as
$=$ $-\frac{1}{2}\sum_{\alpha}\int\frac{\textrm{d}^{2}\delta k\textrm{d}^{2}\delta k^{\prime}}{(2\pi)^{4}}\;|\langle\bm{Q}_{\alpha}+\delta\bm{k}|\bm{Q}_{\alpha}+\delta\bm{k}^{\prime}\rangle|^{2}V_{col}(\delta\bm{k}-\delta\bm{k}^{\prime})n^{0}_{\alpha,\bm{Q}_{\alpha}+\delta\bm{k}}n^{0}_{\alpha,\bm{Q}_{\alpha}+\delta\bm{k}^{\prime}}$ (S1)
$=$ $-\frac{1}{2}\sum_{\alpha\neq\beta}\int\frac{\textrm{d}^{2}\delta k\textrm{d}^{2}\delta k^{\prime}}{(2\pi)^{4}}\;|\langle\bm{Q}_{\alpha}+\delta\bm{k}|\bm{Q}_{\beta}+\delta\bm{k}^{\prime}\rangle|^{2}V_{col}(\delta\bm{k}-\delta\bm{k}^{\prime}+\bm{Q}_{\alpha}-\bm{Q}_{\beta})n^{0}_{\alpha,\bm{Q}_{\alpha}+\delta\bm{k}}n^{0}_{\beta,\bm{Q}_{\beta}+\delta\bm{k}^{\prime}}$ (S2)
where $\delta k\ll Q_{\alpha}$ is measured from the mini-valley center $\bm{Q}_{\alpha}$. Due to the $1/q$ divergence of the Coulomb interaction $V_{col}(q)$ in Eq. (S1), the lowest order term when expanding with respect to $n_{\alpha}$ is $n_{\alpha}^{3/2}$, which is similar to the conventional 2D electron gas. Then up to $n_{h}^{3}$, the intra-valley exchange energy is expanded as
$E_{ex}^{i}\approx-\sum_{\alpha}(A_{1}n_{\alpha}^{3/2}+A_{2}n_{\alpha}^{5/2})+\mathcal{O}(n_{h}^{7/2})$ (S3)
As for the intra-valley term Eq. (S2), since $\delta k\ll Q_{\alpha}$ is satisfied in the low doping regime, and $\delta\bm{k}-\delta\bm{k}^{\prime}+\bm{Q}_{\alpha}-\bm{Q}_{\beta}$ can never be zero, there will be no singularities in the integration and the expansion of $E_{ex}^{x}$ with respect to $n_{\alpha}$ is the conventional Taylor expansion. Then up to $n_{h}^{3}$, the inter-valley exchange energy should have the expansion form
$E_{ex}^{x}\approx-\sum_{\alpha\neq\beta}(B_{1}n_{\alpha}n_{\beta}+B_{2}n_{\alpha}^{2}n_{\beta})+\mathcal{O}(n_{h}^{4})$ (S4)
By fitting the numerical results, the expansion coefficients are gotten and listed in Tab. 1
Table 1: Expansion coefficients of the exchange energy functional fitted from numerical results.
$A_{1}$ $A_{2}$ $B_{1}$ $B_{2}$
$0.429\;\mathrm{eVnm}$ $-16.0\;\mathrm{eVnm^{3}}$ $0.366\;\mathrm{eVnm^{2}}$ $41.1\;\mathrm{eVnm^{4}}$
II The Dynamics Equation of Valley Concentrations
When integrate over the $\mathcal{Q}_{\alpha}$ valley, the Boltzmann equation becomes
$\frac{\partial n_{\alpha}}{\partial t}=$ $-\int_{\mathcal{Q}_{\alpha}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\;\frac{e}{\hbar}\bm{E}\cdot\nabla_{\bm{k}}n_{\bm{k}}+\frac{2\pi n_{imp}}{\hbar}\int_{\mathcal{Q}_{\alpha}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\int\frac{\textrm{d}^{2}k^{\prime}}{(2\pi)^{2}}\;|T_{\bm{k}\bm{k}^{\prime}}|^{2}\delta(\tilde{\varepsilon}_{\bm{k}}-\tilde{\varepsilon}_{\bm{k}^{\prime}})(n_{\bm{k}^{\prime}}-n_{\bm{k}})$
$=$ $-\int_{\mathcal{Q}_{\alpha}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\;\frac{e}{\hbar}\bm{E}\cdot\nabla_{\bm{k}}n_{\bm{k}}+\frac{2\pi n_{imp}|T_{i}|^{2}}{\hbar}\int_{\mathcal{Q}_{\alpha}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\int_{\mathcal{Q}_{\alpha}}\frac{\textrm{d}^{2}k^{\prime}}{(2\pi)^{2}}\;\delta(\tilde{\varepsilon}_{\bm{k}}-\tilde{\varepsilon}_{\bm{k}^{\prime}})(n_{\bm{k}^{\prime}}-n_{\bm{k}})$
$+\frac{2\pi n_{imp}|T_{x}|^{2}}{\hbar}\int_{\mathcal{Q}_{\alpha}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\int_{\bar{\mathcal{Q}}_{\alpha}}\frac{\textrm{d}^{2}k^{\prime}}{(2\pi)^{2}}\;\delta(\tilde{\varepsilon}_{\bm{k}}-\tilde{\varepsilon}_{\bm{k}^{\prime}})(n_{\bm{k}^{\prime}}-n_{\bm{k}})$ (S5)
where $\bar{\mathcal{Q}}_{\alpha}$ means regions exclude $\mathcal{Q}_{\alpha}$. The three terms describe the field driving, intra- and inter-valley scattering processes respectively. The intra-valley term change sign when switching the indices $\bm{k}$ and $\bm{k}^{\prime}$ which means the integral is zero. This reflects the fact that the intra-valley scattering process will not effect the hole concentrations in each valley. To further simplify the inter-valley term, we approximate the distribution $n_{\bm{k}}$ in the integrand by $n^{0}_{\bm{k}}$ and the last term becomes
$\approx$ $\frac{2\pi n_{imp}|T_{x}|^{2}g^{2}}{\hbar}\sum_{\beta\neq\alpha}\int\textrm{d}\varepsilon\int\textrm{d}\varepsilon^{\prime}\;\delta(\varepsilon+\Sigma_{\alpha}-\varepsilon^{\prime}-\Sigma_{\beta})[\Theta(\mu_{\beta}-\varepsilon^{\prime})-\Theta(\mu_{\alpha}-\varepsilon)]$
$=$ $\frac{2\pi n_{imp}|T_{x}|^{2}g^{2}}{\hbar}\sum_{\beta\neq\alpha}\left[\Theta(\Sigma_{\beta}+\mu_{\beta}-\Sigma_{\alpha}-\mu_{\alpha})\min(\mu_{\beta},\Sigma_{\beta}+\mu_{\beta}-\Sigma_{\alpha}-\mu_{\alpha})-\alpha\leftrightarrow\beta\right].$ (S6)
III Dependence of the Critical Field Strength on Hole Concentration and Relaxation Times
we first consider the zero-doping limit $n_{h}\to 0$. At this time, the dynamics equation of valley concentrations could be much simplified and analytically solved. In the zero doping limit, an electrical field in the direction $\pi/3<\theta<\pi$ can only drive holes from the $\mathcal{Q}_{1}$ to the $\mathcal{Q}_{2}$ valley and we could define the driving rate as
$\eta\equiv$ $\lim_{n_{h}\to 0}-\frac{1}{n_{1}}\int_{\mathcal{Q}_{2}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\;\frac{e}{\hbar}\bm{E}\cdot\nabla_{\bm{k}}n_{\bm{k}}$
$=$ $\lim_{n_{h}\to 0}-\frac{1}{n_{1}}\int_{\mathcal{Q}_{2}}\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\;\int_{-\infty}^{0}\textrm{d}t^{\prime}\;\frac{\textrm{e}^{t^{\prime}/\tau_{i}}}{\tau_{i}}\frac{e}{\hbar}\bm{E}\cdot\nabla_{\bm{k}}n^{0}_{1,\bm{k}+e\bm{E}t^{\prime}/\hbar}(n_{1})$
$=$ $-\int_{-\infty}^{0}\textrm{d}t^{\prime}\;\frac{\textrm{e}^{t^{\prime}/\tau_{i}}}{\tau_{i}}\partial_{t^{\prime}}\int_{\mathcal{Q}_{2}}{\textrm{d}^{2}k}\;\delta(\bm{k}+e\bm{E}t^{\prime}/\hbar-\bm{Q}_{1})$
$=$ $\int_{-\infty}^{0}\textrm{d}t^{\prime}\;{\textrm{e}^{t^{\prime}/\tau_{i}}}\int_{0}^{\infty}{\textrm{d}k_{y}}\;\delta(k_{y}+eEt^{\prime}\sin\theta/\hbar+\sqrt{3}Q/2)\int_{-\infty}^{k_{y}/\sqrt{3}}{\textrm{d}k_{x}}\;\delta(k_{x}+eEt^{\prime}\cos\theta/\hbar+Q/2)$
$=$ $\int_{-\infty}^{0}\textrm{d}t^{\prime}\;{\textrm{e}^{t^{\prime}/\tau_{i}}}\int_{0}^{\infty}{\textrm{d}k_{y}}\;\delta(k_{y}+eEt^{\prime}\sin\theta/\hbar+\sqrt{3}Q/2)\Theta(k_{y}/\sqrt{3}+eEt^{\prime}\cos\theta/\hbar+Q/2)$
$=$ $\int_{-\infty}^{0}\textrm{d}t^{\prime}\;{\textrm{e}^{t^{\prime}/\tau_{i}}}\int_{\max(-\sqrt{3}eEt^{\prime}\cos\theta/\hbar-\sqrt{3}Q/2,0)}^{\infty}{\textrm{d}k_{y}}\;\delta(k_{y}+eEt^{\prime}\sin\theta/\hbar+\sqrt{3}Q/2)$
$=$ $\int_{-\infty}^{0}\textrm{d}t^{\prime}\;{\textrm{e}^{t^{\prime}/\tau_{i}}}\Theta[-\max(eEt^{\prime}\sin\theta/\hbar+\sqrt{3}Q/2,eEt^{\prime}(\sin\theta-\sqrt{3}\cos\theta)/\hbar)]$
$=$ $\int_{-\infty}^{-\sqrt{3}\hbar Q/(2eE\sin\theta)}\textrm{d}t^{\prime}\;\textrm{e}^{t^{\prime}/\tau_{i}}$
$=$ $\frac{\textrm{e}^{-\sqrt{3}\hbar Q/(2eE\tau_{i}\sin\theta)}}{\tau_{i}}$ (S7)
In the zero-doping limit, according to Eq. (S3) and (S4), we also have
$\lim_{n_{h}\to 0}\Sigma_{\alpha}=\lim_{n_{h}\to 0}\partial_{n_{\alpha}}E_{ex}=-3A_{1}n_{\alpha}^{1/2}/2+\mathcal{O}(n_{h}^{3/2})$ (S8a)
$\lim_{n_{h}\to 0}\mu_{\alpha}=n_{\alpha}/g+\mathcal{O}(n_{h}^{2})$ (S8b)
Thus the inter-valley scattering term becomes
$\lim_{n_{h}\to 0}\frac{g}{\tau_{x}}\sum_{\beta\neq\alpha}[\Theta(\tilde{\mu}_{\beta}-\tilde{\mu}_{\alpha})\min(\mu_{\beta},\tilde{\mu}_{\beta}-\tilde{\mu}_{\alpha})-\alpha\leftrightarrow\beta]$
$=$ $\lim_{n_{h}\to 0}\frac{g}{\tau_{i}}\sum_{\beta\neq\alpha}[\Theta(\sqrt{n_{\alpha}}-\sqrt{n_{\beta}})\min(n_{\beta}/g,3A_{1}(\sqrt{n_{\alpha}}-\sqrt{n_{\beta}})/2)-\alpha\leftrightarrow\beta]$
$=$ $\frac{1}{\tau_{i}}\sum_{\beta\neq\alpha}\left[\Theta(n_{\alpha}-n_{\beta})n_{\beta}-\alpha\leftrightarrow\beta\right]$ (S9)
In summary, the dynamics equation for $n_{\alpha}$ in the zero-doping limit is finally simplified to
$\dot{n}_{1}=-\eta n_{1}+\frac{1}{\tau_{x}}[\Theta(n_{1}-n_{2})n_{2}-1\leftrightarrow 2]$
$\dot{n}_{2}=\eta n_{1}+\frac{1}{\tau_{x}}[\Theta(n_{2}-n_{1})n_{1}-1\leftrightarrow 2]$
$\dot{p}=\eta(1-p)+\tau_{x}^{-1}\left[\Theta(p)(1-p)-\Theta(-p)(1+p)\right]=(\eta+\tau_{x}^{-1}\mathrm{sgn}\;p)-(\eta+\tau_{x}^{-1})p$ (S10)
The steady polarization is gotten by $\dot{p}^{S}=0$, which gives
$p^{S}=\frac{\eta-\tau_{x}^{-1}}{\eta+\tau_{x}^{-1}},\;p^{S}\leq 0\;$ (S11a)
$p^{S}=1,\;p^{S}>0.$ (S11b)
$\frac{E_{c}}{E^{*}}=\frac{1}{\sin\theta\ln(\tau_{x}/\tau_{i})}$ (S12)
where $E^{*}=\sqrt{3}\hbar Q/(2e\tau_{i})$ is the electrical field unit.
At $\theta=\pi/2$, the critical field strength in the zero doping limit Eq. (S12) is plotted as a function of $\tau_{x}/\tau_{i}$ in Fig. S1(a) by the blue line. Besides, numerically results by solving the Boltzmann equation at finite doping concentrations are also plotted for comparisons. By taking the relaxation times as $\tau_{x}/\tau_{i}=100$ (the gray dashed in Fig. S1(a)), the critical field strength is also plotted as a function of hole doping strength $n_{h}$ in Fig. S1(b).
Figure S1: (a) Critical field for valley flipping as a function of $\tau_{x}/\tau_{i}$ at fixed electrical field direction $\theta=\pi/2$ and the initial state is $\mathcal{Q}_{1}$-polarized. The blue line for the zero doping limit is directly got by Eq. (S12) and the orange and red lines at finite doping concentrations are numerically solved from the Boltzmann equation. (b) The critical field strength as a function of hole concentrations at $\tau_{x}/\tau_{i}=100$ (the dashed gray line in (a)).
IV Symmetry Restrictions on the Angular Pattern of Current Responses
$=$ $M\bm{j}_{N}-\bm{j}_{H}.$ (S13)
$j_{N,\parallel}(2\pi/3-\theta)=j_{N,\parallel}(\theta)$ (S14a)
$j_{N,\perp}(2\pi/3-\theta)=-j_{N,\perp}(\theta)$ (S14b)
$j_{H}(2\pi/3-\theta)=j_{H}(\theta).$ (S14c)
@@ IV Symmetry Restrictions on the Angular Pattern of Current Responses
$j^{(n)}_{N,\parallel}(\theta+\pi)=\hat{r}^{\prime\prime}\cdot\bm{j}^{(n)}(\theta+\pi)=(-1)^{n+1}j^{(n)}_{N,\parallel}(\theta)$ (S16a)
$j^{(n)}_{N,\perp}(\theta+\pi)=\hat{\theta}^{\prime\prime}\cdot\bm{j}^{(n)}(\theta+\pi)=(-1)^{n+1}j^{(n)}_{N,\perp}(\theta)$ (S16b)
$j^{(n)}_{H}(\theta+\pi)=\hat{\theta}^{\prime\prime}\cdot\bm{j}^{(n)}(\theta+\pi)=(-1)^{n+1}j^{(n)}_{H}(\theta)$ (S16c)

### v2

@@ Appendix F Symmetry Restrictions on the Angular Pattern of Current Responses
$j^{(n)}_{N,\parallel}(\theta+\pi)=\hat{r}^{\prime\prime}\cdot\bm{j}^{(n)}(\theta+\pi)=(-1)^{n+1}j^{(n)}_{N,\parallel}(\theta)$ (42a)
$j^{(n)}_{N,\perp}(\theta+\pi)=\hat{\theta}^{\prime\prime}\cdot\bm{j}^{(n)}(\theta+\pi)=(-1)^{n+1}j^{(n)}_{N,\perp}(\theta)$ (42b)
$j^{(n)}_{H}(\theta+\pi)=\hat{\theta}^{\prime\prime}\cdot\bm{j}^{(n)}(\theta+\pi)=(-1)^{n+1}j^{(n)}_{H}(\theta)$ (42c)
Appendix G Angular Pattern of the Second Order Resistivity
In the main text, we calculated the angular pattern of the second order current responses, which is the second order conductivity. However, the quantities directly measured in the experiments are usually resistivity rather than conductivity. To better compare with the experimental results, we will study the second order resistivity in this part. For simplicity, we only focus on the weak field limit where the field driving effect is insignificant and the perturbation theory still works.
For convenience, we will assume the system is polarized in the $\mathcal{Q}_{3}$ valley. Since the Hall current $\bm{j}_{H}$ is two orders smaller than the normal current $\bm{j}_{N}$ as illustrated in Fig. 6, we will ignore the Hall current responses in the following discussion. Then up to second order of $E$, the current is written as
$j_{a}=\sum_{b}\sigma^{(1)}_{ab}E_{b}+\sum_{b,c}\sigma^{(2)}_{abc}E_{b}E_{c},$ (43)
where $a,b,c=x,y$ are spatial indices, $\sigma^{(1)}$ and $\sigma^{(2)}$ are the first and second order conductivity tensors respectively. According to the perturbation theory, these conductivity tensors are calculated as
$\sigma^{(1)}_{ab}=$ $\left(-\frac{e}{\hbar}\right)^{2}\tau_{i}\int\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\;n^{0}_{\bm{k}}\partial_{k_{a}}\partial_{k_{b}}\varepsilon^{0}_{\bm{k}},$ (44a)
$\sigma^{(2)}_{abc}=$ $\left(-\frac{e}{\hbar}\right)^{3}\tau_{i}^{2}\int\frac{\textrm{d}^{2}k}{(2\pi)^{2}}\;n^{0}_{\bm{k}}\partial_{k_{a}}\partial_{k_{b}}\partial_{k_{c}}\varepsilon^{0}_{\bm{k}},$ (44b)
which are symmetric with respect to the spatial indices, i.e., $\sigma^{(1)}_{ab}=\sigma^{(1)}_{ba}$ and $\sigma^{(2)}_{abc}=\sigma^{(2)}_{acb}=\sigma^{(2)}_{bac}$. Besides, due to the restriction of the mirror symmetry with respect to the $x$-axis, we have
$j_{x}(E_{x},-E_{y})=j_{x}(E_{x},E_{y}),$ (45a)
$j_{y}(E_{y},-E_{y})=-j_{y}(E_{x},E_{y}),$ (45b)
which requires $\sigma^{(1)}_{xy}=0$ and $\sigma^{(2)}_{xxy}=\sigma^{(2)}_{yyy}=0$. With the only non-zero components of the conductivity tensors, the relation between the current density and electrical field is written as
$j_{x}=\sigma_{xx}^{(1)}E_{x}+\sigma_{xxx}^{(2)}E_{x}^{2}+\sigma^{(2)}_{xyy}E_{y}^{2}$ (46a)
$j_{y}=\sigma^{(1)}_{yy}E_{y}+2\sigma^{(2)}_{xyy}E_{x}E_{y}.$ (46b)
Similarly, the electrical field $\bm{E}$ could also be written as a series of $\bm{j}$
$E_{x}=\rho^{(1)}_{xx}j_{x}+\rho^{(2)}_{xxx}j_{x}^{2}+\rho^{(2)}_{xyy}j_{y}^{2},$ (47a)
$E_{y}=\rho^{(1)}_{yy}j_{y}+2\rho^{(2)}_{xyy}j_{x}j_{y},$ (47b)
where $\rho^{(1)}$, $\rho^{(2)}$ are the first and second order resistivity tensors. Substitute Eq. (47) into Eq. (46) and keep up to the second order of $j$ we have
$j_{x}=$ $\rho^{(1)}_{xx}\sigma^{(1)}_{xx}j_{x}+\{\rho^{(2)}_{xyy}\sigma^{(1)}_{xx}+[\rho^{(1)}_{yy}]^{2}\sigma^{(2)}_{xyy}\}j_{y}^{2}$
+$+\{\rho^{(2)}_{xxx}\sigma^{(1)}_{xx}+[\rho^{(1)}_{xx}]^{2}\sigma^{(2)}_{xxx}\}j_{x}^{2},$ (48a)
$j_{y}=$ $\rho^{(1)}_{yy}\sigma^{(1)}_{yy}j_{y}+2[\rho^{(1)}_{xx}\rho^{(1)}_{yy}\sigma^{(2)}_{xyy}+\rho^{(2)}_{xyy}\sigma^{(1)}_{yy}]j_{x}j_{y}.$ (48b)
Eq. (48) holds for any $j_{x},j_{y}$, thus the coefficients of the series expansion satisfy the following system of equation
$\rho^{(1)}_{xx}\sigma^{(1)}_{xx}=1,$ (49a)
$\rho^{(2)}_{xyy}\sigma^{(1)}_{xx}+[\rho^{(1)}_{yy}]^{2}\sigma^{(2)}_{xyy}=0,$ (49b)
$\rho^{(2)}_{xxx}\sigma^{(1)}_{xx}+[\rho^{(1)}_{xx}]^{2}\sigma^{(2)}_{xxx}=0,$ (49c)
$\rho^{(1)}_{yy}\sigma^{(1)}_{yy}=1,$ (49d)
$\rho^{(1)}_{xx}\rho^{(1)}_{yy}\sigma^{(2)}_{xyy}+\rho^{(2)}_{xyy}\sigma^{(1)}_{yy}=0.$ (49e)
By solving the previous system of equation, the resistivity tensors are expressed by the conductivity tensors as
$\rho^{(1)}_{xx}=$ $[\sigma^{(1)}_{xx}]^{-1}$ (50a)
$\rho^{(1)}_{yy}=$ $[\sigma^{(1)}_{yy}]^{-1}$ (50b)
$\rho^{(2)}_{xxx}=$ $-\sigma^{(2)}_{xxx}[\sigma^{(1)}_{xx}]^{-3}$ (50c)
$\rho^{(2)}_{xyy}=$ $-\sigma^{(2)}_{xyy}[\sigma^{(1)}_{xx}]^{-1}[\sigma^{(1)}_{yy}]^{-2}$ (50d)
In Ref.[27], the angular resolved second order resistivity in the longitudinal direction is measured. When current $\bm{j}=j(\cos\theta,\sin\theta)$ is applied in the $\theta$ direction, the longitudinal electrical field response is
$E_{\parallel}=$ $E_{x}\cos\theta+E_{y}\sin\theta$
$=$ $j[\rho^{(1)}_{xx}\cos^{2}\theta+\rho^{(1)}_{yy}\sin^{2}\theta]$
$+j^{2}[\rho^{(2)}_{xxx}\cos^{2}\theta+3\rho^{(2)}_{xyy}\sin^{2}\theta]\cos\theta.$ (51)
Thus the second order longitudinal resistivity as a function of theta is
$\rho^{(2)}_{\parallel}=[\rho^{(2)}_{xxx}\cos^{2}\theta+3\rho^{(2)}_{xyy}\sin^{2}\theta]\cos\theta.$ (52)
Define a dimensionless parameter
$\alpha\equiv\frac{3\rho^{(2)}_{xyy}}{\rho^{(2)}_{xxx}}=\frac{3[\sigma^{(1)}_{xx}]^{2}\sigma^{(2)}_{xyy}}{[\sigma^{(1)}_{yy}]^{2}\sigma^{(2)}_{xxx}}.$ (53)
Then, according to Eq. (52), the angular pattern of the second order longitudinal resistivity $\rho^{(2)}_{\parallel}$ is totally determined by $\alpha$. At fixed hold density $n_{h}=0.01\times 10^{12}\mathrm{cm}^{-2}$, the parameter $\alpha$ is calculated and plotted as a function of interlayer bias potential $U$ in Fig. 15(a). In Fig. 15(b-d), the angular patterns of the second order longitudinal resistivity are also plotted at different interlayer potential $U$. When the interlayer potential $U$ is small and $\alpha<1$, the resistivity takes its maximum value at the polarization direction and the angular pattern shown in Fig. 15(b) agree quantitatively with that in Ref. [27](FIG.2.b). With the increase of the interlayer potential $U$, the parameter $\alpha$ also grows. When $\alpha>1$, $\rho^{(2)}_{\parallel}$ has two maximum directions distributed symmetrically on the two sides of the mirror line (when the system is polarized in the $\mathcal{Q}_{3}$ valley, the mirror line is the $x$-axis), and the typical angular pattern is illustrated in Fig. 15(d). At the critical value $\alpha=1$, the angular pattern is also plotted in Fig. 15(c) for reference.
Figure 15: (a) The dimensionless parameter $\alpha$ as a function of interlayer potential $U$. The hole density is fixed at $n_{h}=0.01\times 10^{12}\mathrm{cm}^{-2}$ where the system is in the 1-pocket phase. (b-d) The angular patterns of the second order longitudinal resistivity $\rho^{(2)}_{\parallel}(\theta)$ at different interlayer potential $U$. The system is assumed to be polarized in the $\mathcal{Q}_{3}$ valley.